OpenStudy (anonymous):

Simplify the expression show your work 2-4i ---- 1-2i

jhonyy9 (jhonyy9):

so you need multiplie the numerator and denominator by conjugate of denominator do you know wath will be this conjugate of ?

OpenStudy (anonymous):

No i dont really know how to do any of it

jhonyy9 (jhonyy9):

for you can eliminate the term of i from denominator you need multiplie this by conjugate so you dont have learned it again ?

jhonyy9 (jhonyy9):

look for example if a+ib wann being the complex number so than the conjugate of it will be a-ib this will help you finding the conjugate of denominator from your exercise

OpenStudy (anonymous):

i was taught it but ive never grasped it..but that i is ment for imaginiary

jhonyy9 (jhonyy9):

so than can you solve it now here ?

jhonyy9 (jhonyy9):

the conjugate is 1+2i multiplie the numerator and denominator by and please check what will get ok ?

OpenStudy (anonymous):

ok its still kinda confusing

jhonyy9 (jhonyy9):

2-4i (2-4i)(1+2i) ---- = ----------- = 1-2i (1-2i)(1+2i) can you continue ?

jhonyy9 (jhonyy9):

2+4i-4i+8 ----------- = 1+2i-2i+4 can you continue it now ?